# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 4.1. Variants of SWE-Bench Beyond Lite and Verified
Some other sub-benchmarks have emerged from SWE-Bench Lite and SWE-Bench Verified. For instance, Xia et al. (xia2024agentlessdemystifyingllmbasedsoftware, ) constructed SWE-bench Lite-S by filtering out problematic issues from the full SWE-Bench dataset to enable more rigorous evaluation and comparison—similar in spirit to the goals behind SWE-Bench Verified developed by OpenAI. Ma et al. proposed SWE-bench-Lite-FIX(OLDma2024understandRepoUnderstander, ), an evolution of SWE-Bench Lite where 45 issues out of the 300 from Lite were improved (fixed) by augmenting the issue description.
SWE-Bench Verified has also been forked into smaller subsets. One example is SWEBench-verified-mini [^29] https://github.com/mariushobbhahn/SWEBench-verified-mini, a reduced version containing 50 issues. According to the authors, this subset maintains a similar distribution of performance metrics, test pass rates, and difficulty. Its issues are sourced from only two projects (SWE-Bench has 13). lite_and_verified_solvable [^30] https://github.com/aorwall/moatless-tools?tab=readme-ov-file#run-evaluation contains 84 issues that appear in both the Lite and Verified datasets and have at least one successful submission on SWE-Bench.
@@
Although the SWE-Bench leaderboard includes both the SWE-Bench Lite and SWE-Bench Verified subsets, it does not currently track or evaluate submissions on other emerging forks and derivatives (e.g., Lite-S, verified-mini). Without controlled versioning and official support -such as recognition or integration by the leaderboard maintainers, as happened with Verified- these new subsets or evolutions of existing ones may remain underutilized and have limited influence within the research community.
4.2. Inflated results due to the patch overfitting
An overfitting patch is one that passes all test cases, including the bug-revealing test, but is still incorrect due to weaknesses in the validation oracle, which in practice consists of the test suite (smith2015cure, ). Patch overfitting affects not only early APR systems but also recent state-of-the-art approaches, including Passerine from Google (rondon2025AgentGoogle, ) and AutoCoderRover (Zhang2024AutoCoderRover, ). These works performed manual analyses to identify and discard overfitting patches from their reported results. However, it remains unclear to what extent all submitters SWE-Bench are aware of or account for the overfitting problem. Recent studies of SWE-Bench results, such as (wang2025solvedSWEBench, ), have analyzed patch correctness and found a substantial number of overfitting patches, leading to an average overstatement of resolution rates by 6.2 absolute percentage points. This highlights the need for more rigorous evaluation practices beyond tests to ensure the reliability of leaderboard-reported results.
4.3. Nature of the submitters among SWE-Bench Lite and SWE-Bench Verified
When comparing nature of submitters among SWE-Bench Lite and SWE-Bench Verified, results show an increase in company-oriented submissions (embracing Small, Medium, Large and Large Traded categories), which grow from 44% in SWE-Bench Lite to 63% in SWE-Bench Verified. We hypothesize the following reasons: 1) Since SWE-Bench Verified has been introduced as an extension of SWE-Bench by a leading company in the AI field -OpenAI—, companies may be more inclined to adopt a benchmark validated by an industrial actor, rather than relying solely on a purely academic benchmark such as SWE-Bench Lite. 2) The nature of the benchmark is different. As mentioned on the SWE-Bench website, SWE-Bench Lite is focused on functional bug fixes. Academia has a long-standing tradition in the discipline of Automated Program Repair (legoues2019APRACM, ). For this reason, it is reasonable that SWE-Bench Lite aligns more closely with academic research interests, potentially explaining its higher adoption among researchers compared to other subsets.
@@ 5.1. Empirical studies of SWE-Bench results
Previous work have studied in details the effectiveness of approaches evaluated on SWE-Bench with a deep focus on the generated patches. For example, Meng et al. (meng2024empiricalstudyllmbasedagents, ) conducted an empirical study of the efficiency of 7 repair systems on SWE-bench Lite. In particular, they studied the patches from 4 commercial systems (MarsCode Agent (liu2024marscode, ), Honeycomb, Gru, Alibaba Lingma Agent (ma2025alibabalingmaagent, )) and 3 open-source (AutoCodeRover (Zhang2024AutoCoderRover, ), Agentless + RepoGraph (ouyang2024repographenhancingaisoftware, ) , Agentless (xia2024agentlessdemystifyingllmbasedsoftware, )). All of them are published in the ledearboard and analyzed in our study. They conducted a comprehensive analysis of the performance differences of each system. This research has three main goals: studying repairability and quality, effect on the fault localization capability, and impact of the reproducibility on bugfixing, with a foccus on performance difference of each system. They use Issue Quality Metrics initially defined by Xia et al. (xia2024agentlessdemystifyingllmbasedsoftware, ) to qualify each issue. The study found that issue quality significantly influences the effectiveness of resolution methods, underscoring the importance of crafting clear and comprehensive issue descriptions to improve resolution rates from the outset. They authors also found that existing methods have achieved relatively strong performance in file-level fault localization, but there remains room for improvement in line-level localization tasks. Finally, Reproduction can provide additional information for defect localization when issue information is lacking and help verify the accuracy of generated candidate patches. However, when the issue description is already clear and precise, reproduction may mislead the LLM’s judgment, reducing its focus on the issue description.
@@
Chen et al. (chen2025unveilingpitfallsunderstandingaidriven, ) go beyond evaluating the final patch generated by agents and instead analyze the full trajectories and testing logs from eight agents submitted to SWE-Bench Verified. Their study focuses on Python execution failures and the most frequent error types encountered during the issue-solving phase, examining how these influence the final outcomes. They find that a higher frequency of execution errors is associated with lower-quality patches and increased corrective reasoning efforts, ultimately complicating the repair process.
@@ 5.2. Evaluations on SWE-Bench not included in the leaderboards
Other issue repair approaches have been evaluated on SWE-Bench; however, their results have either not been submitted to the official leaderboard or were not published there at the time of writing this paper.
@@
Yang et al. (yang2025enhancing, ) propose KGCompass, an issue repair approach that incorporates a novel repository-aware knowledge graph (KG) designed to accurately link repository artifacts (such as issues and pull requests) with codebase entities (including files, classes, and functions). This graph is used both to narrow the search space during issue localization and to augment LLMs with relevant contextual information during patch generation. In their evaluation on SWE-Bench Lite, KGCompass achieves a resolution rate of 45.67%.
@@ 6. Conclusion
In this paper, we presented the first comprehensive study of the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards. We analyzed 147 entries corresponding to 67 unique approaches, examining their origins, product characteristics, LLM usage, and architectural design. Our results showed that the majority of high-performing submissions relied on proprietary LLMs, particularly Claude 3.5/3.7, and that both agentic and non-agentic solutions were capable of achieving competitive performance. We also found that most submissions came from industry, including both large tech companies and small startups, as well as independent developers. By systematically analyzing the submissions to the SWE-Bench leaderboards, we uncovered the range of architectural choices and organizational profiles behind them, contributing to a deeper understanding of the current state of the field.

### v2

@@ 4. Discussion
4.1. Patch Overfitting on Program Repair and SWE-Bench
An overfitting patch refers to a fix that passes all available test cases, including the bug-revealing one, yet remains incorrect incorrect because of weaknesses in the validation oracle, which in practice consists of the test suite. The software engineering community has a long history of studying patch overfitting, beginning with the seminal works by Smith et al.(smith2015cure, ), who coined the term overfitting patch, and Qi et al.(qi2015analysis, ) et al., who introduced the concept of plausible patches, patches that pass all test cases yet incorrect.
Nevertheless, patch overfitting affects not only early APR systems but also recent state-of-the-art approaches, including Passerine from Google (rondon2025AgentGoogle, ), Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) and AutoCoderRover (Zhang2024AutoCoderRover, ). In these works, the authors conducted manual analyses to identify and exclude overfitting patches from their reported results on SWE-Bench, following a practice commonly adopted in earlier APR research evaluated on other benchmarks, such as Defects4J (e.g., (martinez2017automatic, ; Liu2020:ontheefficiency, )). However, it remains unclear to what extent all submitters SWE-Bench are aware of or account for the overfitting problem. Recent studies of SWE-Bench results, such as (wang2025solvedSWEBench, ), have analyzed patch correctness and found a substantial number of overfitting patches, leading to an average overstatement of resolution rates by 6.2 absolute percentage points.
Academic or research-oriented submissions -specially those that have been involved in the software engineering community (e.g. (xia2024agentlessdemystifyingllmbasedsoftware, ; ruan2024specrovercodeintentextraction, ; Zhang2024AutoCoderRover, ; rondon2025AgentGoogle, ; ma2025alibabalingmaagent, )) are likely to be aware of the quality standards expected by the research community, for example, validating patch correctness beyond plausibility, often through manual inspection or even eventually with automated techniques such as DDR (Ye2021:DDR, ). In their papers, these authors typically report the ratio of bugs repaired with correct patches as well as those with merely plausible ones. However, we have not observe such analyzes in articles that targets the artificial intelligence community (e.g. (yang2024sweagentagent, ; li2025patchpilot, ; zainullina2025guidedsearchstrategies, )). This highlights a misalignment between the software engineering and AI communities in evaluation standards: while the former emphasizes semantic correctness, the latter often relies solely on test-suite-based validation, potentially overlooking incorrect yet test-passing patches. Additionally, some submissions —particularly from industry— may not be fully aware of the underlying issues associated with test-suite-based evaluation, such as patch overfitting. These observations highlight the need for the software engineering community to more effectively disseminate established evaluation practices and engage in cross-community dialogue to help align standards more broadly.
The SWE-Bench framework and leaderboards provide a valuable foundation for reproducible experimentation and progress tracking in automated issue repair. However, it currently lacks mechanisms for deeper correctness validation beyond passing test cases. As a result, it remains unclear whether submitters—particularly those outside academia—perform additional analyses to verify the semantic correctness of their patches. This underscores the need for more rigorous evaluation practices to complement test-based validation. Future work could enhance the SWE-Bench platform by incorporating optional or standardized procedures for post-submission correctness assessment, helping to ensure the reliability of reported results.
@@ 4.2. Variants of SWE-Bench Beyond Lite and Verified
Some other sub-benchmarks have emerged from SWE-Bench Lite and SWE-Bench Verified. For instance, Xia et al. (xia2024agentlessdemystifyingllmbasedsoftware, ) constructed SWE-bench Lite-S by filtering out problematic issues from the full SWE-Bench dataset to enable more rigorous evaluation and comparison—similar in spirit to the goals behind SWE-Bench Verified developed by OpenAI. Ma et al. proposed SWE-bench-Lite-FIX (OLDma2024understandRepoUnderstander, ), an evolution of SWE-Bench Lite where 45 issues out of the 300 from Lite were improved (fixed) by augmenting the issue description. SWE-Bench Verified has also been forked into smaller subsets. One example is SWEBench-verified-mini [^27] https://github.com/mariushobbhahn/SWEBench-verified-mini, a reduced version containing 50 issues. According to the authors, this subset maintains a similar distribution of performance metrics, test pass rates, and difficulty. Its issues are sourced from only two projects (SWE-Bench has 13). Another subset, named lite_and_verified_solvable  [^28] https://github.com/aorwall/moatless-tools?tab=readme-ov-file#run-evaluation, contains 84 issues that appear in both the Lite and Verified datasets and have at least one successful submission on SWE-Bench.
@@
Although the SWE-Bench leaderboard includes both the SWE-Bench Lite and SWE-Bench Verified subsets, it does not currently track or evaluate submissions on other emerging forks and derivatives (e.g., Lite-S, verified-mini). Without controlled versioning and official support -such as recognition or integration by the leaderboard maintainers, as happened with SWE-Bench Verified- these new subsets or evolutions of existing ones may remain underutilized and have limited influence within the research community.
4.3. The debate: Single-Agent or Multiple- Agent architecture?
In our study, we found that the majority of submissions correspond to the Agentic-based approach, and these type of solutions can be Single Agent or Multiple Agents. From observing the results from both Lite and Verified leaderboards, we cannot determine that one architecture achieves better results than the other. However, different players have publicly endorsed one approach over the other.
For instance, Cognition, creators of one of the Devin pioneering AI software engineer agent, argued against building multi-agent systems. They consider such systems fragile because agents often lack access to a shared global context, which is crucial for making coherent, high-quality decisions. At the same time, researchers have also focused on failures of multi-agent systems. For instance, Cemri et al.(cemri2025multiagentllmsystemsfail, ) examined the trajectories of six multi-agent systems, including HyperAgent (phan2024hyperagent, ), a system with results submitted to the SWE-Bench Lite dashboard. Based on 200 task executions, the authors identified 14 distinct failure modes, grouped into three overarching categories: 1) Specification issues (failures stemming from poor system design or ambiguous prompt specifications), 2) Inter-agent misalignment (failures caused by breakdowns in coordination and communication between agents during execution), and 3) Task verification (failures due to inadequate validation or premature task termination). While not all failures were directly attributable to the multi-agent architecture, the most prevalent ones fell under the inter-agent misalignment category —highlighting issues that are inherently linked to the challenges of coordinating multiple agents.
In contrast, just one day later of that post from Cognition, Anthropic published a post outlining lessons learned from building their research system based on a multi-agent architecture. [^29] https://www.anthropic.com/engineering/built-multi-agent-research-system They argue that, for their use case, sub-agents are beneficial because they operate in parallel with separate context windows, enabling better compression. Additionally, sub-agents support separation of concerns by using distinct tools, prompts, and exploration paths —reducing path dependency and allowing for more thorough and independent investigations. The authors also acknowledge that certain domains, such as most coding tasks, are not well suited for multi-agent systems with current technologies, as they typically require all agents to share the same context or involve complex interdependencies between agents. This reveals a partial alignment between the positions of both companies, despite their differing architectural choices.
We also found discussions about this architectural issue on the meta-data from SWE-Bench entries. For example, the entry Lingxi focuses on a multi-agent approach, which offers a way to split the repair workflow into compact, purpose-built agents that each receive only the information they need. That offered an option to single agent pipelines where keep the entire dialogue in one context inevitably suffer from context dilution, where the prompt is dominated by earlier discussion tokens. On the contrary, OpenHands fosters a single-agent architecture (neubig2024openhandsmultiagent, ) characterized by a single a) LLM, b) action space, and c) prompting technique. The authors identify several challenges with multi-agent architectures, including: 1) designing an effective agentic structure, where each agent has clear responsibilities and dedicated tools; 2) preserving context across multiple agents; and 3) maintaining the agents’ codebases and prompt.
Other submitters to SWE-Bench have evolved the architecture of their approaches across submissions to address issues arising from their initial choice of single-agent or multi-agent design. For instance, nFactorial has submitted four entries to SWE-Bench Verified, each featuring a distinct system architecture. Their initial submission followed a non-agentic design. In the second submission, they transitioned to a multi-agent architecture incorporating specialized agents such as the Localizer, Fixer, and Code Editor. However, in the third submission, they merged the Localizer and Fixer into a single agent. According to the authors, this change was motivated by the observation that passing context between separate agents sometimes resulted in the loss of critical information. According to the authors of Warp (71% on Verified), they initially explored multi-agent approaches —including dedicated testing, planning, and reasoning agents— but ultimately found that a single primary agent yielded the most consistent and reliable results.
In conclusion, there is no one-size-fits-all solution, no single architecture prevails, as submitters navigate trade-offs between the modularity of multi-agent systems and the simplicity and context integrity of single-agent designs. Foundational frameworks and practical toolsets like MAST, proposed by Cemri et al. (cemri2025multiagentllmsystemsfail, ), offer valuable means for diagnosing and improving both single- and multi-agent systems. These tools can help identify and mitigate the types of failures and weaknesses highlighted in the mentioned articles and studies.
4.4. Known Limitations of SWE-Bench
Despite the success of SWE-Bench —adopted by researchers across diverse communities, including Software Engineering and Artificial Intelligence, and having a notable impact on industry, as discussed in Section 3.1.2— it still shares some limitations common to other bug and issue benchmarks. These limitations have even been acknowledged and discussed by several submitters. For example, Zencoder [^30] https://zencoder.ai/blog/demystifying-swe-bench summarizes the following. (1) Data Contamination: Many issues in SWE-Bench were created before the training cutoff dates of several LLMs. This raises the risk of data leakage, where models may have been exposed to issue descriptions or solutions during training, potentially inflating performance scores. To address this, researchers have developed infrastructures to ensure freshness and isolation of benchmark data. For instance, Zhang et al. (zhang2025swebenchgoeslive, ) proposed SWE-bench-Live, an automated curation pipeline that streamlines the full process, from instance creation to environment setup, allowing the benchmark to be continuously updated with recent data. (2) Lack of Real-World Representation: SWE-Bench focuses exclusively on Python repositories, which may limit its ability to represent the full diversity of programming languages and software engineering tasks encountered in real-world development. For example, Zeng et al. (zeng2025skyworkswe, ) introduced Skywork-SWE, a dataset comprising 10,169 real-world Python task instances. Their evaluation using the OpenHands scaffold showed significantly lower resolve rates than those observed on SWE-Bench —e.g., 20.23% with Gemini-2.5-Pro, 18.54% with GPT-4.1, and 15.94% with o3-mini— suggesting potential overestimation of model performance on SWE-Bench. (3) Benchmark Specificity: Scores on SWE-Bench may reflect a model’s performance on known or curated datasets rather than its ability to generalize to novel, diverse, or noisy real-world scenarios.
4.5. Approaching Saturation: Observations from SWE-Bench Trends
In July 2025, submissions to SWE-Bench achieved up to 75% precision—an increase of more than 20 percentage points compared to the previous year. This progress coincides with the release of increasingly capable LLMs. As depicted in Figure 9(b), all systems exceeding 70% precision relied on state-of-the-art Claude 4 models, either alone or in combination with others. If such trends persist, upcoming LLM releases could surpass 80% or even 90%, suggesting that SWE-Bench may be approaching a saturation point, as previously observed with benchmarks like HumanEval for the code generation task. [^31] https://paperswithcode.com/sota/code-generation-on-humaneval
4.6. The importance of Open-source frameworks for experimentation
SWE-Bench serves as an excellent example of the value of open-source tools in enabling rapid experimentation and research progress. Beyond the benchmark infrastructure is open-source, one of the key factors contributing to its early success was the availability of open-source implementations of several initial submissions—particularly the tools or scaffolds used in approaches such as Agentless, AutoCodeRover, SWE-Agent, Moatless, OpenHands, and PatchPilot. This openness has significantly accelerated the pace of research by providing a variety of accessible frameworks on which researchers, developers, and industry practitioners can build, extend, and evaluate new ideas. As discussed in Section 3.2.1, numerous contributions have extended and adapted Agentless, demonstrating the benefits of having a reusable and modifiable open-source foundation. A notable example of this openness is the SIMA submission (27.67% on Lite), where a single developer experimentally evaluated the ideas proposed in a paper on multi-agent architectures (li2024MoreAgentsneed, ) by implementing them on top of the open-source Moatless tool. This highlights how open-source frameworks lower the barrier to entry and empower individual contributors to engage in impactful research and benchmarking.
Surprisingly, some submitters from industry have also embraced open-source practices. For instance, the small company Augment Code, in addition to offering commercial products (e.g., development plugins), has open-sourced the agent used in their SWE-Bench evaluation. [^32] Augment Code Agent: https://github.com/augmentcode/augment-swebench-agent/tree/main Notably, this agent held the highest precision on the Verified leaderboard (65.4%) as of March 2025. This open-source strategy not only enhances the transparency and reproducibility of their results but also contributes to the broader advancement of the community—including other industrial players—as demonstrated by its influence on ByteDance’s submission TRAE, which adopted Augment Code’s open-sourced patch selector.
5. Threats to Validity
External Validity.
In this study, we focused on two SWE-Bench leaderboards, which we selected due to the SWE-bench’s substantial impact on both academia and industry, as demonstrated both by the companies involved (most industry big players) and the scientific impact, e.g. with Jimenez et al.’s seminal paper reaching over 780 citations on Google Scholar as of July 2025. Other benchmarks may be equally representative of the broader issue-fixing landscape, but we do not claim that our findings can be applied to them.
Internal Validity.
We followed a systematic procedure to collect information from the leaderboard submissions, enabling us to characterize both the submitters and the approaches. However, there is a risk that we may have missed documents describing some approaches, resulting in certain submissions being left uncharacterized. To mitigate this, we extended our search beyond the official leaderboard metadata, consulting multiple sources including Google searches, LinkedIn profiles, and scientific publications indexed on Google Scholar and arXiv. Furthermore, the entire information-gathering process was repeated twice to ensure coverage and consistency.
Construct Validity.
Our classification was based on a published taxonomy (liu2024SurveyAgentsLLM4SE, ) and three architectural dimensions (workflow authoring, control flow autonomy and presence of agents). While these constructs were designed to capture core aspects of LLM-based repair systems, applying them consistently was challenging because many submissions lacked detailed or standardized descriptions of their solutions. To mitigate potential misclassification, we used explicit categories for unknown cases and cross-validated all annotations among the authors.
Some potentially important dimensions and variables were not analyzed in this study. First, we did not account for the number of parameters or the specific minor version of each LLM. This decision was made to simplify the analysis and focus on identifying broader trends across model families and major releases, rather than fine-grained differences between specific versions. Second, we did not evaluate the monetary cost associated with the different submissions, despite some submitters reporting this information. While cost is a relevant factor for assessing the usability and practicality of an approach, especially given that most of the submissions rely or proprietary (paid) models, it was excluded from our analysis for practical reasons. Token prices have generally decreased over time, making cost comparisons across submissions potentially misleading without proper normalization. Performing such normalization would require detailed, often unavailable, information from the submissions (e.g., price per token at the time of experimentation, total tokens consumed). Due to these limitations in cost reporting and data availability, we chose to exclude monetary cost from our study.
Conclusion Validity.
Some submissions, especially those without formal publications and described only in blog posts or informal channels, lacked essential details (e.g. related to the solution architecture - single, multiple or no agent, autonomy) or contained ambiguous information (e.g., the LLMs used in experimentation by the submitters). To avoid misclassification, we introduced special categories to explicitly denote unknown or missing information. Additionally, cross-validation of the classification decisions was conducted among the authors to further reduce the risk of bias or error.
@@ 6.1. Empirical studies of SWE-Bench results
Previous work have studied in details the effectiveness of approaches evaluated on SWE-Bench with a deep focus on the generated patches. For example, Meng et al. (meng2024empiricalstudyllmbasedagents, ) conducted an empirical study of the efficiency of 7 repair systems on SWE-bench Lite. In particular, they studied the patches from 4 commercial systems (MarsCode Agent (liu2024marscode, ), Honeycomb, Gru, Alibaba Lingma Agent (ma2025alibabalingmaagent, )) and 3 open-source (AutoCodeRover (Zhang2024AutoCoderRover, ), Agentless + RepoGraph (ouyang2024repographenhancingaisoftware, ) , Agentless (xia2024agentlessdemystifyingllmbasedsoftware, )). All of them are published in the ledearboard and analyzed in our study. They conducted a comprehensive analysis of the performance differences of each system. This research has three main goals: studying repairability and quality, effect on the fault localization capability, and impact of the reproducibility on bugfixing, with a focus on performance difference of each system. They use Issue Quality Metrics initially defined by Xia et al. (xia2024agentlessdemystifyingllmbasedsoftware, ) to qualify each issue. The study found that issue quality significantly influences the effectiveness of resolution methods, underscoring the importance of crafting clear and comprehensive issue descriptions to improve resolution rates from the outset. They authors also found that existing methods have achieved relatively strong performance in file-level fault localization, but there remains room for improvement in line-level localization tasks. Finally, Reproduction can provide additional information for defect localization when issue information is lacking and help verify the accuracy of generated candidate patches. However, when the issue description is already clear and precise, reproduction may mislead the LLM’s judgment, reducing its focus on the issue description.
@@
Chen et al. (chen2025unveilingpitfallsunderstandingaidriven, ) go beyond evaluating the final patch generated by agents and instead analyze the full trajectories and testing logs from eight agents submitted to SWE-Bench Verified. Their study focuses on Python execution failures and the most frequent error types encountered during the issue-solving phase, examining how these influence the final outcomes. They find that a higher frequency of execution errors is associated with lower-quality patches and increased corrective reasoning efforts, ultimately complicating the repair process. Bouzenia et Pradel (bouzenia2025understanding, ) studied the trayectories of three agents, two with results submitted to SWE-Bench (AutoCodeRover, OpenHands) and one RepairAgent (bouzenia2024RepairAgent, ), not previously evaluated in (bouzenia2024RepairAgent, ) but in Defects4J (Just2014Defects4J, ). They categorize the actions (Explore, Locate, Search, Generate fix, etc.) and then mined frequent sequences of categorized actions in agent trajectories to identify recurrent decision-making pattern. They found for instance, that successful trajectories (that conduct to correct patches) balance exploration, explanation, and validation steps, while unsuccessful ones exhibit repetitive, non-adaptive action cycles.
@@ 6.2. Evaluations on SWE-Bench not included in the leaderboards
Further issue repair approaches have been evaluated on SWE-Bench, although their results were either not submitted to or not published on the leaderboards at the time of writing.
@@ 7. Conclusion
In this paper, we presented the first comprehensive study of the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards. We analyzed 178 entries corresponding to 80 unique approaches from 71, examining their origins, product characteristics, LLM usage, and architectural design of the approaches. Our results showed that the majority of high-performing submissions relied on proprietary LLMs, particularly Claude 4, and that both agentic and non-agentic solutions were capable of achieving competitive performance. We also found that most submissions came from industry, including both large tech companies and small startups, as well as independent developers. By systematically analyzing the submissions to the SWE-Bench leaderboards, we uncovered the range of architectural choices and organizational profiles behind them, contributing to a deeper understanding of the current state of the field.